A cubical block of mass m and edge 'a' slides down a rough inclined plane of inclination θ with uniform velocity. Find the torque of the normal force acting on the block, about its centre.

Grade:12th pass

1 Answers

Arun
5 years ago
The force mg acting on the body has two components mg sin θ and mg cos θ and the body will exert a normal reaction. Let R = Since R and mg cos θ pass through the centre of the cube, there will be no torque due to R and mg cos θ. The only torque will be produced by mg sin θ. i = F × r (r = a/2) (a = ages of the cube) i = mg sin θ × a/2 = 1/2 mg a sin θ.